OpenStudy (hba):

Equation of circle with center at (root(2),-3root(3) ) and radius 2 root (2)

OpenStudy (hba):

No one answers here...Thank you.

OpenStudy (hba):

I am solving it .

OpenStudy (anonymous):

oh shush, mi trying to remember the circle equation

OpenStudy (anonymous):

You have done this earlier @hba

OpenStudy (anonymous):

\[(x-x_0)^2+y-y_0)^2=r^2\]

OpenStudy (hba):

x^2+y^2+2root(2)x-6root(3)y+2root(2)=0 @waterineyes :D Right or wrong @hartnn Yes ?

OpenStudy (hba):

My radius is wrong lol :D

OpenStudy (anonymous):

Yes you have forgot to square it I guess..

OpenStudy (anonymous):

Also It will come to RHS..

OpenStudy (hba):

Wait i will do it again thanks :)

OpenStudy (anonymous):

Equation is also wrong.. Everything is wrong. Do it again. ha ha ha..

OpenStudy (hba):

I know lol ha ha ha

OpenStudy (anonymous):

Take your time..

OpenStudy (anonymous):

Sorry, take your time but solve it within 30 seconds.. Ha ha ha..

OpenStudy (hba):

I know that x^2+y^2-2root(2) x +6 root(3) y +21=0

OpenStudy (hba):

Here you go @waterineyes

OpenStudy (anonymous):

Where I go???

OpenStudy (anonymous):

Looking good to me..

OpenStudy (anonymous):

2 + 27 = 8 implies 21 on LHS..

OpenStudy (hba):

err....Right or wrong ?

OpenStudy (anonymous):

Looking good means it is right.. Looking bad means it is absolutely right.. Okay?? Ha ha ha..

OpenStudy (hba):

Looks bad to me :)

OpenStudy (anonymous):

Well Done, you are right this time and now I am serious...